# AMPLIFY:Attention-based Mixup for Performance Improvement and Label Smoothing in Transformer

### v2

@@ 3. Method
    (5) $\Lambda=\left\{\lambda_{i}=\text{WS}\left(\text{Beta}\left(\alpha,\alpha\right)\right)\right\}_{i=1}^{n}=\left\{\lambda_{1},\lambda_{2},\ldots,\lambda_{n}\right\},~{}\lambda_{\text{max}}=\text{max}\left(\Lambda\right)$
@@
    (6) $\begin{split}x_{i}^{\text{mix}}&=\lambda_{\text{max}}\cdot x_{i}^{o}+\left(1-\lambda_{\text{max}}\right)\cdot x_{i}^{s}\\ y_{i}^{\text{mix}}&=\lambda_{\text{max}}\cdot y_{i}^{o}+\left(1-\lambda_{\text{max}}\right)\cdot y_{i}^{s}\\ \textbf{M}_{B}^{\text{mix}}=\left\{\left\langle x_{i}^{\text{mix}},y_{i}^{\text{mix}}\right\rangle\right\}_{i=1}^{l}&=\left\{\left\langle x_{1}^{\text{mix}},y_{1}^{\text{mix}}\right\rangle,\left\langle x_{2}^{\text{mix}},y_{2}^{\text{mix}}\right\rangle,\ldots,\left\langle x_{l}^{\text{mix}},y_{l}^{\text{mix}}\right\rangle\right\}\end{split}$
@@
  • Step 6: Calculate the loss value $\mathcal{L}_{\text{mix}}$ based on the predicted results of the model. Standard Mixup uses two common methods to calculate the loss value. One method calculates the cross-entropy loss value by taking the $logits$ output by the text classification head and computing it with the ground truth in both the original order and the shuffled order, and then weighting the two loss values and adding them together as the total loss. The other method calculates the cross-entropy loss value between the $logits$ and the mixed pseudo-labels. Essentially, the main difference between these two methods is the order in which the weighting and cross-entropy loss calculations are performed. The first method calculates the cross-entropy loss value of the results first, and then performs weighting and summation, while the second method first performs weighting and summation of the results before calculating the corresponding cross-entropy loss value. In the AMPLIFY algorithm, according to the experimental results of (Yoon et al., 2021), we adopt the more effective first method to calculate the loss value $\mathcal{L}_{\text{mix}}$. The specific process can be represented by the following equations:
@@
    (7) $logits=\mathcal{F}_{N}\left(\textbf{M}_{B}^{\text{mix}}\right),~{}index_{o}=I\left(\textbf{M}_{B}^{o}\right)$
@@
    (8) $\begin{split}gt_{o}=\text{LABEL}\left(\textbf{M}_{B}^{o},index_{o}\right)&,~{}gt_{s}=\text{LABEL}\left(\textbf{M}_{B}^{o},index_{R}\right)\\ \mathcal{L}_{\text{mix}}\Leftarrow\lambda_{\text{max}}&\cdot CrossEntropy\left(logits,gt_{o}\right)+\\ \left(1-\lambda_{\text{max}}\right)&\cdot CrossEntropy\left(logits,gt_{s}\right)\end{split}$
@@
Input: The pre-trained text classification model based on Transformer architecture $T_{\text{pre}}$
@@
Output: The Logits output by the text classification head $logits$
@@
       Use equation 2 to obtain the shuffled mini-batch $\textbf{M}_{B}^{s}$ and its index $index_{R}$. foreach multi-head attention layer $\in T_{\text{pre}}$ do
@@
4. Experiments
4.1. Benchmark Datasets and Models
When designing our experiment, considering the representativeness of the model and its relevance to text classification tasks, we chose the BERT-base-uncased model  (Devlin et al., 2018) from the HuggingFace Transformers library as the backbone network among many pre-trained models based on the Transformer architecture. we conducted experiments on seven benchmark datasets including :
  • MRPC: The MRPC dataset is a binary classification task dataset used for determining whether sentences have similar meanings. It consists of sentence pairs sourced from the internet, where the pairs are annotated as either "yes" or "no" to indicate whether they have the same semantics. The dataset comprises 3669 sentence pairs, with each pair having a binary label (Wang et al., 2018).
  • SST-2: The SST-2 dataset is a binary classification task dataset used for sentiment classification, specifically classifying the sentiment polarity of sentences as either positive or negative. It contains 6921 samples, with each sample having a binary label (Wang et al., 2018).
  • SST-5: The SST-5 dataset is a fine-grained sentiment classification task dataset used for sentiment classification. It divides sentences into five emotional categories, namely very negative, negative, neutral, positive, and very positive. The dataset comprises 8545 samples, with each sample having a five-class sentiment label (Socher et al., 2013).
  • TREC: The TREC dataset is a dataset used for question classification tasks, classifying questions into coarse-grained and fine-grained categories. It encompasses diverse types of questions, including those related to person names, locations, numbers, and more. The dataset consists of 5453 question samples, with six coarse-grained labels and 50 fine-grained labels (Li and Roth, 2002).
  • Yelp-5: The YELP-5 dataset is a sentiment classification task dataset used for sentiment analysis. It comprises user reviews sourced from the Yelp website. The task involves categorizing the reviews into five sentiment categories: very negative, negative, neutral, positive, and very positive. The dataset consists of 650,000 user review samples, with each sample having a five-class sentiment label (Yelp, 2014).
  • IMDB: The IMDB dataset is a sentiment classification task dataset used for sentiment analysis. It consists of sentences extracted from movie reviews. The objective of the task is to determine whether the reviews are positive or negative in sentiment. The dataset comprises 25,001 movie review samples, with each sample having a binary sentiment label (Maas et al., 2011).
@@ 4.2. Baselines
  • No Mixup: Relying solely on the predictive ability of the backbone network without using Mixup technology  (Devlin et al., 2018).
  • EmbedMix: First, the zero-padding technique is used to pad all text sequences in the training set to the same length. After completing word embedding processing, pairs of sequences are randomly combined. Then, Mixup operations are performed separately on the two vectorized samples in the sequence pairs and their corresponding classification labels to obtain the augmented sample and its pseudo-label. Therefore, the Mixup operation of this method occurs in the word embedding stage of the text preprocessing, involving only semantic features in the word vector space  (Guo et al., 2019a).
  • SentenceMixup: First, the encoder of the text model is used to process all samples in the training set to obtain the corresponding sentence-level sequence encoding. Then, Mixup operations are performed separately on the two randomly selected sequence encodings and their labels to obtain the feature sequence after linear interpolation and its pseudo-label. Finally, the mixed result is fed to the softmax layer at the end of the network. Therefore, the Mixup operation of this method occurs in the prediction stage of text processing, and the entire feature aggregation process only involves the hidden layers within the classification head  (Guo et al., 2019a).
  • TMix: First, two sequences $x_{a}$ and $x_{b}$ are randomly selected from the training set and processed by the text model $T$. Then, a random hidden layer is selected from the hidden layers of $T$, and the output $x_{a}^{f}$ and $x_{b}^{f}$ of $x_{a}$ and $x_{b}$ in that hidden layer are extracted. Then, Mixup operations are performed separately on these two feature sequences $x_{a}^{f}$ and $x_{b}^{f}$ and their labels, obtaining mixed feature sequence after linear interpolation and its corresponding pseudo-label, which are then fed to the subsequent hidden layer. Therefore, the Mixup operation of this method occurs in the feature extraction stage of text processing, involving only a specific hidden layer in the model  (Chen et al., 2020).
@@
(9) $\lambda_{t}=\text{WS}\left(\text{Beta}\left(\alpha,\alpha\right)\right),~{}\lambda=\text{max}\left(\lambda_{t},1-\lambda_{t}\right)$
@@
4.3. Experimental Settings
In all experiments, AdamW  (Loshchilov and Hutter, 2017) is chosen as the optimizer for training, using cosine learning rate  (Shazeer and Stern, 2018), warmup step accounting for 10% of the total training steps, initial learning rate of 2e-5, EPS of 1e-8, weight decay coefficient of 0.01, batch size of 32, maximum sequence length of 256, maximum epochs of 15 for fine-tuning the pre-trained model, and early stop patience of 5. To ensure the consistency and effectiveness of the experimental process, the construction method of all neural network models involved comes from HuggingFace Transformers  (Wolf et al., 2019), and all experiments are completed on the same NVIDIA RTX A6000 GPU based on the same configuration file parameters under the Pytorch Lightning framework. Each experiment uses 3 different random seeds and reports the mean and variance of the results.
5. Results
5.1. Overall Results
@@
Table 2. Comparison experimental results of different Mixup methods on seven benchmark datasets. All values in the table are the average accuracy (%) and variance of the model after running three times with three different random seeds  (Dror et al., 2018).
@@
Figure 3. The figure presents a heatmap of the attention output matrices comparing our method with EmbedMix, SentenceMix, TMix, and the baseline. Each element in the matrix represents the p-value obtained from conducting a t-test between our method and the other methods. Different color distributions are used to differentiate the magnitude of p-values.
@@
What caught our attention is that on the Yelp-5 dataset with 560,000 training samples, almost all Mixup methods failed to provide a net performance gain for the baseline model. Observing the experimental process and results, We believe this is because pre-trained models like BERT generally perform well when facing large-scale datasets, as they are usually pre-trained on massive corpora of textual data, allowing them to learn more language-level knowledge and patterns. Therefore, they have already achieved excellent performance on datasets such as Yelp-5, which have relatively clear classification features, leaving limited room for Mixup to improve their performance. On the other hand, Mixup methods essentially augment the samples and their labels by using the existing feature information in the same dataset, which is different from standard data augmentation strategies and does not introduce out-of-domain feature information. Therefore, they cannot significantly improve the model’s performance or seriously impair it. When the dataset is large, the baseline model can already understand the text features well. In this case, using Mixup methods may not have a significant effect. However, when the dataset is small and the model’s overfitting problem is severe, like many other data augmentation methods, Mixup can often have a significant effect, helping the model improve its generalization ability and robustness in the face of sample uncertainty  (Sun et al., 2020).
@@
For example, the MRPC dataset has only 4500 samples, and using the Mixup method on it can lead to performance gains, especially for AMPLIFY, which performs multiple Mixup operations during the model’s prediction process, effectively adding multiple mild random perturbations to the feature sequence, resulting in more significant performance gains. Additionally, when using the Mixup method to augment feature sequences, the mixed sequences must differ significantly from the original ones to improve the model’s generalization performance. If the number of sequences per class in the dataset is relatively balanced, the distribution of differences between the mixed samples will also be more uniform, making it difficult for the Mixup method to bring further performance gains, as demonstrated in the study by  (Yoon et al., 2021). From this perspective, the sample sizes of various categories in Yelp-5 are very similar, and the feature distribution of the data is relatively balanced. Therefore, the differences between the mixed sequences are not significant enough, which renders the Mixup method ineffective. Furthermore, on SST-2, the AMPLIFY method resulted in negative gains in model performance. After analyzing the reasons, we found that SST-2 is a dataset used for sentiment binary classification tasks, and it contains text samples of audience comments on movies or annotations of audience sentiment on movies. Therefore, these comment texts vary greatly in length, and after padding, many text sequences contain a large number of meaningless placeholders. Therefore, when performing Mixup operations on feature sequences of short and long texts, it may mix placeholders with sentiment information, which weakens or even submerges the classification features, thereby affecting the model’s prediction results.
@@
Figure 3 illustrates a heatmap of the p-values obtained from t-tests comparing our method with other methods for the attention output matrices in each dataset we computed. Each cell represents the t-test p-value between our method and other methods. It is worth noting that in this heatmap, the p-values between our method and all other Mixup methods are less than 0.05, indicating significant differences. These statistically significant differences should not be overlooked. They suggest that our method exhibits notable distinctions from other Mixup methods in generating attention matrices, further emphasizing the superiority of our approach. By analyzing these significant differences, we can gain a deeper understanding of the improvements our method brings to attention output and quantify their statistical significance. Firstly, we can discuss the quality of the attention matrices. Through the attention analysis from Figure 6 to Figure 9, our method demonstrates superior performance in terms of stability, consistency, and accuracy compared to other methods. This implies that our method is better able to capture crucial features and allocate attention weights more accurately. Furthermore, the impact of these significant differences on practical tasks is also noteworthy. Experimental results indicate that our method exhibits better performance in specific tasks. Such practical performance disparities further strengthen the advantages resulting from the improvements in attention output brought about by our method. Additionally, a notable aspect of our method among Mixup approaches lies in how it leverages attention mechanisms to enhance the effectiveness of Mixup. This uniqueness positions our method as more advantageous in attention output compared to other methods and further enhances the overall performance.
@@
To further validate the potential issue of Mixup propagating noise or outlier features from the original samples to the augmented samples, leading to over-sensitivity of the model, as well as to demonstrate the effectiveness of our method in improving generalization and denoising, we conducted additional experiments on five different datasets by randomly deleting or swapping original data at proportions of 5%, 10%, 15%, and 20%. These experiments aimed to investigate the performance of different Mixup methods in the presence of noisy datasets. The results in Figure 4 demonstrate that our method achieved the highest accuracy in almost all datasets. We attribute this success to the introduction of attention mechanisms in our method, which allows the model to focus on important features and adaptively adjust attention weights during the Mixup process. This adaptive adjustment of attention weights enables our method to handle variations introduced by randomly deleting or swapping data samples more effectively. By dynamically allocating attention weights, the model can effectively handle missing or swapped data, resulting in more accurate predictions. Furthermore, the performance improvement of our method, AMPLIFY, can also be attributed to its utilization of inherent dependencies and correlations within the data through attention mechanisms. By attending to relevant regions or features, AMPLIFY can effectively propagate useful information among samples, thereby enhancing generalization and accuracy.
@@
Table 3. The effectiveness of using AMPLIFY on DistilBERT.
Figure 4. Results of Four Mixup Methods on SST-2 , SST-5 , MRPC , TREC-coarse and TREC-fine Training Sets with Random Deleted and Swapped of Data at Various Proportions of 5%, 10%, 15%, and 20%
5.2. Variance
@@
5.3. Visualization of Experimental Results
Figure 5. The cross-entropy loss values for four Mixup methods, EmbedMix, TMix, AMPLIFY, and SentenceMix, on the MRPC dataset in the first 12k training iterations.
Figure 6. The influence of the fourth MHA layer of the model on the same text sequence after undergoing AMPLIFY and EmbedMix operations, respectively. The left side of the figure represents the word being updated, while the right side represents the word being processed. The lines in the figure represent the semantic correlations between words, and the color depth reflects the weight of attention obtained from the correlation. The text sequence consists of two sentences "the rabbit quickly hopped" and "the turtle slowly crawled", with [SEP] being a special token used to separate the two sentences and [CLS] being a special token used to classify the text sequence.
Figure 7. The influence of the fourth MHA layer of the model, after undergoing AMPLIFY and EmbedMix operations, respectively, on a specific word "rabbit".
Figure 8. The neuron view of how the fourth MHA layer of the model calculates attention weights based on query and key vectors after being mixed by two Mixup methods, AMPLIFY and EmbedMix, respectively. The positive values are represented in blue, while the negative values are represented in orange. The depth of the color indicates the weight, and the lines represent the attention between words. The input text sequence consists of two identical sentences, "The animal didn’t cross the street because it was too tired".
Figure 9. The semantic relationship graph between words corresponding to the output of the first MHA layer, with or without the AMPLIFY operation. The color depth of the nodes represents different attention weights, and the currently focused word is highlighted. The input text sequence consists of two identical sentences, "A coming-of-age film that avoids the cartoonish clichés and sneering humor of the genre as it provides a fresh view of an old type".
@@ 5.3. Visualization of Experimental Results
Figure 7 shows the effect of the two Mixup operations on the attention mechanism when applied to a specific word in the same sentence. Clearly, AMPLIFY does not have a negative impact on the output of MHA, the separator is successfully recognized, and closely related words still have a high attention weight, such as the two words "rabbit" and "hopped" which have the highest attention weight. In contrast, EmbedMix has a negative impact on the output of MHA, making it difficult to select the correct word and assign appropriate attention weights, resulting in the separator not playing its role and the attention being scattered.

### v3

@@ Introduction
    $\Lambda=\left\{\lambda_{i}=\text{WS}\left(\text{Beta}\left(\alpha,\alpha\right)\right)\right\}_{i=1}^{n}=\left\{\lambda_{1},\lambda_{2},\ldots,\lambda_{n}\right\},\leavevmode\nobreak\ \lambda_{\text{max}}=\text{max}\left(\Lambda\right)$ (5)
@@
    $\begin{split}x_{i}^{\text{mix}}&=\lambda_{\text{max}}\cdot x_{i}^{o}+\left(1-\lambda_{\text{max}}\right)\cdot x_{i}^{s}\\ y_{i}^{\text{mix}}&=\lambda_{\text{max}}\cdot y_{i}^{o}+\left(1-\lambda_{\text{max}}\right)\cdot y_{i}^{s}\\ \textbf{M}_{B}^{\text{mix}}=\left\{\left\langle x_{i}^{\text{mix}},y_{i}^{\text{mix}}\right\rangle\right\}_{i=1}^{l}&=\left\{\left\langle x_{1}^{\text{mix}},y_{1}^{\text{mix}}\right\rangle,\left\langle x_{2}^{\text{mix}},y_{2}^{\text{mix}}\right\rangle,\ldots,\left\langle x_{l}^{\text{mix}},y_{l}^{\text{mix}}\right\rangle\right\}\end{split}$ (6)
@@
  • Step 6: Calculate the loss value $\mathcal{L}_{\text{mix}}$ based on the predicted results of the model. Standard Mixup uses two common methods to calculate the loss value. One method calculates the cross-entropy loss value by taking the $logits$ output by the text classification head and computing it with the ground truth in both the original order and the shuffled order, and then weighting the two loss values and adding them together as the total loss. The other method calculates the cross-entropy loss value between the $logits$ and the mixed pseudo-labels. Essentially, the main difference between these two methods is the order in which the weighting and cross-entropy loss calculations are performed. The first method calculates the cross-entropy loss value of the results first, and then performs weighting and summation, while the second method first performs weighting and summation of the results before calculating the corresponding cross-entropy loss value. In the AMPLIFY algorithm, according to the experimental results of Yoon et al., (2021), we adopt the more effective first method to calculate the loss value $\mathcal{L}_{\text{mix}}$. The specific process can be represented by the following equations:
@@
    $logits=\mathcal{F}_{N}\left(\textbf{M}_{B}^{\text{mix}}\right),\leavevmode\nobreak\ index_{o}=I\left(\textbf{M}_{B}^{o}\right)$ (7)
@@
    $\begin{split}gt_{o}=\text{LABEL}\left(\textbf{M}_{B}^{o},index_{o}\right)&,\leavevmode\nobreak\ gt_{s}=\text{LABEL}\left(\textbf{M}_{B}^{o},index_{R}\right)\\ \mathcal{L}_{\text{mix}}\Leftarrow\lambda_{\text{max}}&\cdot CrossEntropy\left(logits,gt_{o}\right)+\\ \left(1-\lambda_{\text{max}}\right)&\cdot CrossEntropy\left(logits,gt_{s}\right)\end{split}$ (8)
@@
Input:
The pre-trained text classification model based on Transformer architecture $T_{\text{pre}}$
@@
Output:
The Logits output by the text classification head $logits$
@@
      Use equation 2 to obtain the shuffled mini-batch $\textbf{M}_{B}^{s}$ and its index $index_{R}$.
      foreach multi-head attention layer $\in T_{\text{pre}}$ do
@@
Experiments
Benchmark Datasets and Models
When designing our experiment, considering the representativeness of the model and its relevance to text classification tasks, we chose the BERT-base-uncased model  Devlin et al., (2018) from the HuggingFace Transformers library as the backbone network among many pre-trained models based on the Transformer architecture. we conducted experiments on seven benchmark datasets including :
  • MRPC [^1] https://huggingface.co/datasets/SetFit/mrpc: The MRPC dataset is a binary classification task dataset used for determining whether sentences have similar meanings. It consists of sentence pairs sourced from the internet, where the pairs are annotated as either ”yes” or ”no” to indicate whether they have the same semantics. The dataset comprises 3669 sentence pairs, with each pair having a binary label Wang et al., (2018).
  • SST-$2$https://huggingface.co/datasets/SetFit/sst2: The SST-2 dataset is a binary classification task dataset used for sentiment classification, specifically classifying the sentiment polarity of sentences as either positive or negative. It contains 6921 samples, with each sample having a binary label Wang et al., (2018).
  • SST-5333https://huggingface.co/datasets/SetFit/sst5: The SST-5 dataset is a fine-grained sentiment classification task dataset used for sentiment classification. It divides sentences into five emotional categories, namely very negative, negative, neutral, positive, and very positive. The dataset comprises 8545 samples, with each sample having a five-class sentiment label Socher et al., (2013).
  • TREC [^4] https://www.kaggle.com/datasets/thedevastator/the-trec-question-classification-dataset-a-longi: The TREC dataset is a dataset used for question classification tasks, classifying questions into coarse-grained and fine-grained categories. It encompasses diverse types of questions, including those related to person names, locations, numbers, and more. The dataset consists of 5453 question samples, with six coarse-grained labels and 50 fine-grained labels Li and Roth, (2002).
  • Yelp-$5$http://www.yelp.com/datasetchallenge: The YELP-5 dataset is a sentiment classification task dataset used for sentiment analysis. It comprises user reviews sourced from the Yelp website. The task involves categorizing the reviews into five sentiment categories: very negative, negative, neutral, positive, and very positive. The dataset consists of 650,000 user review samples, with each sample having a five-class sentiment label Yelp, (2014).
  • IMDB [^6] https://www.kaggle.com/datasets/ashirwadsangwan/imdb-dataset: The IMDB dataset is a sentiment classification task dataset used for sentiment analysis. It consists of sentences extracted from movie reviews. The objective of the task is to determine whether the reviews are positive or negative in sentiment. The dataset comprises 25,001 movie review samples, with each sample having a binary sentiment label Maas et al., (2011).
@@
  • No Mixup: Relying solely on the predictive ability of the backbone network without using Mixup technology  Devlin et al., (2018).
  • EmbedMix: First, the zero-padding technique is used to pad all text sequences in the training set to the same length. After completing word embedding processing, pairs of sequences are randomly combined. Then, Mixup operations are performed separately on the two vectorized samples in the sequence pairs and their corresponding classification labels to obtain the augmented sample and its pseudo-label. Therefore, the Mixup operation of this method occurs in the word embedding stage of the text preprocessing, involving only semantic features in the word vector space  Guo et al., 2019a .
  • SentenceMixup: First, the encoder of the text model is used to process all samples in the training set to obtain the corresponding sentence-level sequence encoding. Then, Mixup operations are performed separately on the two randomly selected sequence encodings and their labels to obtain the feature sequence after linear interpolation and its pseudo-label. Finally, the mixed result is fed to the softmax layer at the end of the network. Therefore, the Mixup operation of this method occurs in the prediction stage of text processing, and the entire feature aggregation process only involves the hidden layers within the classification head  Guo et al., 2019a .
  • TMix: First, two sequences $x_{a}$ and $x_{b}$ are randomly selected from the training set and processed by the text model $T$. Then, a random hidden layer is selected from the hidden layers of $T$, and the output $x_{a}^{f}$ and $x_{b}^{f}$ of $x_{a}$ and $x_{b}$ in that hidden layer are extracted. Then, Mixup operations are performed separately on these two feature sequences $x_{a}^{f}$ and $x_{b}^{f}$ and their labels, obtaining mixed feature sequence after linear interpolation and its corresponding pseudo-label, which are then fed to the subsequent hidden layer. Therefore, the Mixup operation of this method occurs in the feature extraction stage of text processing, involving only a specific hidden layer in the model  Chen et al., (2020).
@@
$\lambda_{t}=\text{WS}\left(\text{Beta}\left(\alpha,\alpha\right)\right),\leavevmode\nobreak\ \lambda=\text{max}\left(\lambda_{t},1-\lambda_{t}\right)$ (9)
@@
Experimental Settings
In all experiments, AdamW  Loshchilov and Hutter, (2017) is chosen as the optimizer for training, using cosine learning rate  Shazeer and Stern, (2018), warmup step accounting for 10% of the total training steps, initial learning rate of 2e-5, EPS of 1e-8, weight decay coefficient of 0.01, batch size of 32, maximum sequence length of 256, maximum epochs of 15 for fine-tuning the pre-trained model, and early stop patience of 5. To ensure the consistency and effectiveness of the experimental process, the construction method of all neural network models involved comes from HuggingFace Transformers  Wolf et al., (2019), and all experiments are completed on the same NVIDIA RTX A6000 GPU based on the same configuration file parameters under the Pytorch Lightning framework. Each experiment uses 3 different random seeds and reports the mean and variance of the results.
Results
Overall Results
@@ Results
Table 2: Comparison experimental results of different Mixup methods on seven benchmark datasets. All values in the table are the average accuracy (%) and variance of the model after running three times with three different random seeds  Dror et al., (2018).
@@
Figure 3: The figure presents a heatmap of the attention output matrices comparing our method with EmbedMix, SentenceMix, TMix, and the baseline. Each element in the matrix represents the p-value obtained from conducting a t-test between our method and the other methods. Different color distributions are used to differentiate the magnitude of p-values.
@@
What caught our attention is that on the Yelp-5 dataset with 560,000 training samples, almost all Mixup methods failed to provide a net performance gain for the baseline model. Observing the experimental process and results, We believe this is because pre-trained models like BERT generally perform well when facing large-scale datasets, as they are usually pre-trained on massive corpora of textual data, allowing them to learn more language-level knowledge and patterns. Therefore, they have already achieved excellent performance on datasets such as Yelp-5, which have relatively clear classification features, leaving limited room for Mixup to improve their performance. On the other hand, Mixup methods essentially augment the samples and their labels by using the existing feature information in the same dataset, which is different from standard data augmentation strategies and does not introduce out-of-domain feature information. Therefore, they cannot significantly improve the model’s performance or seriously impair it. When the dataset is large, the baseline model can already understand the text features well. In this case, using Mixup methods may not have a significant effect. However, when the dataset is small and the model’s overfitting problem is severe, like many other data augmentation methods, Mixup can often have a significant effect, helping the model improve its generalization ability and robustness in the face of sample uncertainty  Sun et al., (2020).
@@
For example, the MRPC dataset has only 4500 samples, and using the Mixup method on it can lead to performance gains, especially for AMPLIFY, which performs multiple Mixup operations during the model’s prediction process, effectively adding multiple mild random perturbations to the feature sequence, resulting in more significant performance gains. Additionally, when using the Mixup method to augment feature sequences, the mixed sequences must differ significantly from the original ones to improve the model’s generalization performance. If the number of sequences per class in the dataset is relatively balanced, the distribution of differences between the mixed samples will also be more uniform, making it difficult for the Mixup method to bring further performance gains, as demonstrated in the study by  Yoon et al., (2021). From this perspective, the sample sizes of various categories in Yelp-5 are very similar, and the feature distribution of the data is relatively balanced. Therefore, the differences between the mixed sequences are not significant enough, which renders the Mixup method ineffective. Furthermore, on SST-2, the AMPLIFY method resulted in negative gains in model performance. After analyzing the reasons, we found that SST-2 is a dataset used for sentiment binary classification tasks, and it contains text samples of audience comments on movies or annotations of audience sentiment on movies. Therefore, these comment texts vary greatly in length, and after padding, many text sequences contain a large number of meaningless placeholders. Therefore, when performing Mixup operations on feature sequences of short and long texts, it may mix placeholders with sentiment information, which weakens or even submerges the classification features, thereby affecting the model’s prediction results.
@@
Figure 3 illustrates a heatmap of the p-values obtained from t-tests comparing our method with other methods for the attention output matrices in each dataset we computed. Each cell represents the t-test p-value between our method and other methods. It is worth noting that in this heatmap, the p-values between our method and all other Mixup methods are less than 0.05, indicating significant differences. These statistically significant differences should not be overlooked. They suggest that our method exhibits notable distinctions from other Mixup methods in generating attention matrices, further emphasizing the superiority of our approach. By analyzing these significant differences, we can gain a deeper understanding of the improvements our method brings to attention output and quantify their statistical significance. Firstly, we can discuss the quality of the attention matrices. Through the attention analysis, our method demonstrates superior performance in terms of stability, consistency, and accuracy compared to other methods. This implies that our method is better able to capture crucial features and allocate attention weights more accurately. Furthermore, the impact of these significant differences on practical tasks is also noteworthy. Experimental results indicate that our method exhibits better performance in specific tasks. Such practical performance disparities further strengthen the advantages resulting from the improvements in attention output brought about by our method. Additionally, a notable aspect of our method among Mixup approaches lies in how it leverages attention mechanisms to enhance the effectiveness of Mixup. This uniqueness positions our method as more advantageous in attention output compared to other methods and further enhances the overall performance.
@@
To further validate the potential issue of Mixup propagating noise or outlier features from the original samples to the augmented samples, leading to over-sensitivity of the model, as well as to demonstrate the effectiveness of our method in improving generalization and denoising, we conducted additional experiments on five different datasets by randomly deleting or swapping original data at proportions of 5%, 10%, 15%, and 20%. These experiments aimed to investigate the performance of different Mixup methods in the presence of noisy datasets. The results in Table 4 and 5 demonstrate that our method achieved the highest accuracy in almost all datasets. We attribute this success to the introduction of attention mechanisms in our method, which allows the model to focus on important features and adaptively adjust attention weights during the Mixup process. This adaptive adjustment of attention weights enables our method to handle variations introduced by randomly deleting or swapping data samples more effectively. By dynamically allocating attention weights, the model can effectively handle missing or swapped data, resulting in more accurate predictions. Furthermore, the performance improvement of our method, AMPLIFY, can also be attributed to its utilization of inherent dependencies and correlations within the data through attention mechanisms. By attending to relevant regions or features, AMPLIFY can effectively propagate useful information among samples, thereby enhancing generalization and accuracy.
@@
Table 3: The effectiveness of using AMPLIFY on DistilBERT.
Dataset Percentage No Mixup EmbedMix SentenceMix TMix Ours
MRPC 5% 68.52 67.30 68.41 68.46 73.04
10% 67.83 67.25 66.84 67.77 72.87
15% 67.30 67.30 68.23 67.19 72.64
20% 67.25 68.99 68.93 68.81 71.65
SST-5 5% 39.55 37.87 40.63 39.19 44.30
10% 37.96 36.02 37.56 36.61 43.12
15% 39.23 37.60 40.50 39.32 43.26
20% 34.71 35.43 38.64 36.61 42.62
SST-2 5% 85.67 86.60 86.82 86.99 87.48
10% 86.38 86.00 86.44 86.44 86.66
15% 84.68 83.96 85.28 85.61 87.48
20% 83.53 83.03 82.87 83.47 85.61
IMDB 5% 91.49 91.51 91.56 91.64 91.78
10% 91.55 91.72 91.61 91.67 91.89
15% 91.67 91.64 91.57 91.42 91.81
20% 91.45 91.50 91.55 91.60 91.65
TREC-coarse 5% 95.80 96.60 96.40 96.20 96.80
10% 96.80 96.40 96.20 95.20 97.00
15% 96.40 95.80 95.40 96.40 96.60
20% 96.00 95.80 96.10 95.90 96.40
TREC-fine 5% 87.00 85.80 87.60 86.20 87.80
10% 86.40 86.20 87.00 87.20 87.60
15% 85.20 85.40 84.40 84.40 85.60
20% 84.60 84.20 84.20 84.80 85.80
Table 4: Results of Four Mixup Methods on SST-2 , SST-5 , MRPC , TREC-coarse and TREC-fine Training Sets with Random Deleted of Data at Various Proportions of 5%, 10%, 15%, and 20%
Dataset Percentage No Mixup EmbedMix SentenceMix TMix Ours
MRPC 5% 67.88 66.90 68.75 67.01 72.93
10% 67.65 64.35 67.19 66.67 71.30
15% 66.20 65.16 63.01 59.94 70.72
20% 65.10 65.04 63.13 66.90 67.30
SST-5 5% 40.65 40.95 40.95 41.95 43.67
10% 36.86 35.88 34.98 37.06 44.25
15% 38.16 38.55 37.96 38.14 41.36
20% 36.60 22.04 35.20 38.64 38.96
SST-2 5% 81.76 85.94 87.31 86.11 87.59
10% 84.97 85.17 84.46 85.34 87.77
15% 84.88 83.42 84.68 85.50 87.59
20% 85.39 82.59 82.92 82.04 87.15
IMDB 5% 91.42 91.46 91.40 91.56 91.74
10% 91.43 91.43 90.96 90.79 91.56
15% 90.25 89.82 90.36 90.28 90.56
20% 86.51 88.80 88.48 88.35 89.08
TREC-coarse 5% 97.00 96.40 97.00 96.60 97.80
10% 95.80 96.60 96.20 96.20 96.80
15% 94.60 95.60 95.60 95.60 95.80
20% 96.20 96.00 95.60 95.40 96.80
TREC-fine 5% 86.60 86.40 86.40 87.40 87.60
10% 85.80 84.80 86.00 84.80 86.40
15% 84.60 84.00 84.40 84.60 85.80
20% 84.80 85.00 86.00 83.60 85.80
Table 5: Results of Four Mixup Methods on SST-2 , SST-5 , MRPC , TREC-coarse and TREC-fine Training Sets with Random Swapped of Data at Various Proportions of 5%, 10%, 15%, and 20%
Variance
@@
Visualization of Experimental Results
Figure 4: The cross-entropy loss values for four Mixup methods, EmbedMix, TMix, AMPLIFY, and SentenceMix, on the MRPC dataset in the first 12k training iterations.
Figure 5: The influence of the fourth MHA layer of the model on the same text sequence after undergoing AMPLIFY and EmbedMix operations, respectively. The left side of the figure represents the word being updated, while the right side represents the word being processed. The lines in the figure represent the semantic correlations between words, and the color depth reflects the weight of attention obtained from the correlation. The text sequence consists of two sentences ”the rabbit quickly hopped” and ”the turtle slowly crawled”, with [SEP] being a special token used to separate the two sentences and [CLS] being a special token used to classify the text sequence.
Figure 6: The influence of the fourth MHA layer of the model, after undergoing AMPLIFY and EmbedMix operations, respectively, on a specific word ”rabbit”.
Figure 7: The neuron view of how the fourth MHA layer of the model calculates attention weights based on query and key vectors after being mixed by two Mixup methods, AMPLIFY and EmbedMix, respectively. The positive values are represented in blue, while the negative values are represented in orange. The depth of the color indicates the weight, and the lines represent the attention between words. The input text sequence consists of two identical sentences, ”The animal didn’t cross the street because it was too tired”.
Figure 8: The semantic relationship graph between words corresponding to the output of the first MHA layer, with or without the AMPLIFY operation. The color depth of the nodes represents different attention weights, and the currently focused word is highlighted. The input text sequence consists of two identical sentences, ”A coming-of-age film that avoids the cartoonish clichés and sneering humor of the genre as it provides a fresh view of an old type”.
@@
Figure 6 shows the effect of the two Mixup operations on the attention mechanism when applied to a specific word in the same sentence. Clearly, AMPLIFY does not have a negative impact on the output of MHA, the separator is successfully recognized, and closely related words still have a high attention weight, such as the two words ”rabbit” and ”hopped” which have the highest attention weight. In contrast, EmbedMix has a negative impact on the output of MHA, making it difficult to select the correct word and assign appropriate attention weights, resulting in the separator not playing its role and the attention being scattered.